OpenStudy (anonymous):

Which two values of x roots are polynomial below 5x^2+5x-1

OpenStudy (anonymous):

\[A. x=\frac{-8\pm \sqrt{28} }{ 6 }\]

OpenStudy (anonymous):

B.\[x=\frac{ 5\pm \sqrt{5} }{ 10}\]

OpenStudy (anonymous):

C.\[x=\frac{ 5\pm \sqrt{45} }{ 20 }\]

hero (hero):

Use the quad formula

OpenStudy (anonymous):

the ? doesnt mean solve using the quadratic formula right because i solved and i got something different

OpenStudy (anonymous):

i did but my answer doesnt match none of this

OpenStudy (anonymous):

oh snaps damn srewed up i got B xD

OpenStudy (anonymous):

i mean c

hero (hero):

So you figured it out now?

OpenStudy (anonymous):

nevermin i got \[x=\frac{ -5\pm \sqrt{45} }{ 10}\]

OpenStudy (mathstudent55):

\(5x^2+5x-1 = 0\) \( x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) Here, a = 5, b = 5, and c = -1, so \( x = \dfrac{-5 \pm \sqrt{5^2 - 4(5)(-1)}}{2(5)} \) \( x = \dfrac{-5 \pm \sqrt{25 + 20}}{10} \) \( x = \dfrac{-5 \pm \sqrt{45}}{10} \) \( x = \dfrac{-5 \pm 3\sqrt{5}}{10} \) None of the answers are correct.

OpenStudy (anonymous):

wait the 1 is +

OpenStudy (anonymous):

the answer is B sorry guys

OpenStudy (mathstudent55):

B has +5 instead of -5. Also, B has sqrt(5), not sqrt(45). Look art my step before the last one and compare it to B.

OpenStudy (anonymous):

it is b because my equation is suppose to look like this 5x^2-5x+1

OpenStudy (mathstudent55):

In other words, we just wasted 20 minutes solving the wrong problem and arguing about how the solution choices are all wrong.

hero (hero):

Technically, even though that is true to an extent, he was still able to figure out his own mistake and figure out the correct answer to the correct problem, so it wasn't a total waste.